OpenStudy (anonymous):

How can you algebraically find the limit of the following as n approaches infinite?

OpenStudy (anonymous):

sqrt(n+1)/sqrt(4n+1)

OpenStudy (anonymous):

\[\sqrt{n+1}/\sqrt{4n+1}=\sqrt{n+1}*\sqrt{4n+1}/4n+1\]